Raymond is as twice as old as Rona now. Twenty years ago Raymond was six times as old as Rona. What are there present ages?
Answer by Maths68(1474)

Raymond is as twice as old as Rona now. Twenty years ago Raymond was six times as old as Rona. What are there present ages?
Let
Present age of Raymond = r
Present age of Rona = a
Raymond is as twice as old as Rona now
r=2a...........(1)
Twenty years ago means we have to subtract 20 from their present age
r-20=a-20
Then
Raymond was six times as old as Rona
r-20=6(a-20)
r-20=6a-120
r-6a=-120+20
r-6a=-100
Put the value of r from (1) to above equation
r-6a=-100
2a-6a=-100
-4a=-100
-4a/-4=-100/-4
a=25
Put the value of a in (1)
r=2a
r=2(25)
r=50
Present age of Raymond = r = 50 years old
Present age of Rona = a = 25 years old
Check
Twenty years ago Raymond was six times as old as Rona
Raymond=50-20=30
Rona=25-20=5
Raymond age = 6 * Rona age
30=6*5
30=30